OpenStudy (anonymous):

A ball is thrown upward with an initial velocity of 96 ft/sec from a height 640 ft. It's height h,in feet, after tseconds is given by h(t)= -16t^2 +96t+640. After how long will the ball reach the ground? Tha ball will reach the ground in ___seconds

OpenStudy (anonymous):

When it reaches the ground, \[\ h(t)=0\] Therefore \[ -16t^2 +96t+640=0\]

OpenStudy (anonymous):

It is a quadratic equation, so there will be 2 solutions (which should be obvious to you, as there are 2 points when it touches the ground, before it was thrown and after it was thrown).

OpenStudy (anonymous):

With the question probably asking you for the later time.

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

Do you understand this?

OpenStudy (anonymous):

Nope. All I got was h(t) = -16t^2 + 96t + 640 = 0 ==> t^2 - 6t - 40 = 0

OpenStudy (anonymous):

You understand why -16t^2 + 96t + 640 = 0, though? I'll help with the rest.

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

t^2 - 6t - 40 is the same as (t-10)(t+4)

OpenStudy (anonymous):

So for (t-10)(t+4) to be =0, t =10 or =-4